Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below: Sales (13,200 units × $30 per unit) $ 396,000 Variable expenses 237,600 Contribution margin 158,400 Fixed expenses 176,400 Net operating loss $ (18,000 ) 5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $60,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,700 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,700)?
Answer #1

1)

Total

Perunit

Sales (13,200 units × $30 per unit)

396,000

30

100%

Variable expenses

237,600

18

60%

Contribution margin

158,400

12

40%

Total

Per unit

Sales (20,700 units × $30 per unit)

621,000

30

100%

Variable expenses (20,700 * (18 - 3)

310,500

15

50%

Contribution margin

310,500

15

50%

 

Break-even point in units = (Fixed expenses + 60,000) / Unit contribution margin = (176,400 60,000) / 15 = 15,760 units

(in sales) = (176,400 + 60,000) / 0.50 = $472,800

2)

NOT AUTOMATED

AUTOMATED

Sales (20,700 units × $30 per unit)

621,000

30

100%

621,000

30

100%

Variable expenses (20,700 * $18 per unit)

372,600

18

60%

310,500

15

50%

Contribution margin

248,400

12

40%

310,500

15

50%

Less: Fixed expenses

176,400

236,400

Net operating income

72,000

74,100